) A hollow cylindrical metallic pipe is open at both the ends and its external diameter is 12 cm. If
the length of the pipe is 70 cm and the thickness of the metal used is 1 cm, find the volume of the
metal used for making the pipe

Grade:12th pass

1 Answers

Harshit Singh
2 years ago
Dear Student

volume of cylinder = pi*r2*h

volume of metal used to make pipe = pi*h* (r12-r22)
V= 3.14 * 70 *(72-62)= 220*13= 2860


Thanks